How many grams of carbon dioxide (CO2) will be created from the incomplete combustion of 40 g of octane (C8H18)?
2 C8H18 + 25O2→ 16CO2 + 18H2O
1
Expert's answer
2020-10-26T14:50:53-0400

M(CO2)=12+2×16=44(g/mol),

M(C8H18)=8×12+18×1=114(g/mol).

Mass of CO2 may be found from the proportion:

x=40×16×44/2×114=123.5(g).